A melting point test of n = 10 samples of a binder used in manufacturing a rocket propellant resulted in a sample mean of 154.4 °F. Assume that the melting point is normally distributed with σ = 1.5°F. Is it possible to conclude at 1% significance that the melting point is not 155°F? What is the value of zcalc ? Please report your answer in 3 decimal places.

b: What is the p-value for p(z<-1.3)? Please report your answer in 3 decimal places.
Answer #1

a)

Zcal = - / / sqrt(n)

= 154.4 - 155 / (1.5 / sqrt(10) )

= -1.265

b)

p-value = P ( Z < z)

= P( Z < -1.3)

= 1 - P( Z < 1.3)

= 1 - 0.9032

= 0.097